OpenStudy (anonymous):

If 2x + 7 = 3 and Bx-10 =- 2 and x is the same on the previous questions, what is B?

OpenStudy (lalaly):

find x from first one 2x=3-7 2x=-4 x=-2 now substitute it in the second and find B

OpenStudy (anonymous):

So B is 8 or am I wrong?

OpenStudy (lalaly):

-2B-10=-2 -2B=8 B=-4

OpenStudy (anonymous):

Damn, I suck at math but thanks for the reply.

OpenStudy (lalaly):

you dont suck, u just need to keep practicing.. and your welcome:)